OpenStudy (zyberg):

Find the roots of x^3 - x + 2x^2 - 2 = 0

OpenStudy (amorfide):

you want to try values of x such as 1,2,3,4,5,6 until it gives 0 x=1 1-1+2-2 hence we get 3-3=0 so 1 is a solution so we have x=1, x-1=0 so (x-1)(ax^2+bx+c)=x^3-x+2x^2-2

OpenStudy (amorfide):

so when we multiply this out we get ax^3+bx^2+cx-ax^2-bx-c=x^3-x+2x^2-2 so we have ax^3+bx^2-ax^2+cx-c=x^3-x+2x^2-2 compare coefficients so for x^3 a=1 for x^2 b-a=2 for constants -c=-2 c=2 so b-a=2, a=1, b-1=2, b=3 so we have (x-1)(x^2+3x+2)=x^3-x+2x^2-2